OpenStudy (anonymous):

Given: 2x – 3(x – 5 + x) = 2x – 3 Prove: x = 3

OpenStudy (anonymous):

Mathematical Statement Justification 2x – 3(x - 5 + x) = 2x - 3 Given 2x – 3(x + x – 5) = 2x – 3 Commutative Property of Addition 2x – 3(2x – 5) = 2x – 3 Combine Like Terms 2x – 6x + 15 = 2x – 3 (2x – 6x) + 15 = 2x – 3 – 4x + 15 = 2x – 3 – 6x + 15 = – 3 Subtraction Property of Equality – 6x = – 18 Subtraction Property of Equality x = 3 Division Property of Equality What are the missing justification steps, in order? Answer Subtraction Property of Equality, Commutative Property of Addition, Addition Property of Equality Distributive Property, Commutative Property of Addition, Combine Like Terms Distributive Property, Associative Property of Addition, Combine Like Terms Subtraction Property of Equality, Associative Property of Addition, Addition Property of Equality

OpenStudy (anonymous):

flvs?

OpenStudy (anonymous):

yesD:

OpenStudy (anonymous):

holy crap u a flvs student too?

OpenStudy (anonymous):

me too D: what mod are you in?

OpenStudy (anonymous):

yes haha & im in mod 4 on proofs

OpenStudy (anonymous):

i'm about to finish first segment in algebra 1

OpenStudy (anonymous):

this is like segment 2 geometry .

OpenStudy (anonymous):

btw, i think the answer is property of addition

OpenStudy (anonymous):

me tooo! except im skipping around this is really annoying -.- but w,ee

OpenStudy (anonymous):

<freshman

OpenStudy (anonymous):

sophhh> im skipping around too haha i do the tests first and look up the answers on google then attempt the projects ahah. & the answer choices are What are the missing justification steps, in order? Answer Subtraction Property of Equality, Commutative Property of Addition, Addition Property of Equality Distributive Property, Commutative Property of Addition, Combine Like Terms Distributive Property, Associative Property of Addition, Combine Like Terms Subtraction Property of Equality, Associative Property of Addition, Addition Property of Equality

OpenStudy (anonymous):

ok lets see...

OpenStudy (anonymous):

i think its the last 1

OpenStudy (anonymous):

oh darn it wasnt the last one D: it was the only question i missed.

OpenStudy (anonymous):

oh awesome we could possibly have the same teacher haha

OpenStudy (anonymous):

who do you have?

OpenStudy (anonymous):

i have ms.drisen they pretty much do the same exact thing . Have you done a collab for segment 2 yet?

OpenStudy (anonymous):

nope i actuallt need to do that very soon lol

OpenStudy (anonymous):

haha same here we should collab on the module 9 project thingy

OpenStudy (anonymous):

we should if you wanna get it done soon cause i end school jun4 so i need to get it done before then

OpenStudy (anonymous):

mine ends around that time too . Wanna add me on facebook so we could talk on there?

OpenStudy (anonymous):

yeah sure just send me a message with your link on here or via message

OpenStudy (anonymous):

i know this is a late reply but some else might need help with this Distributive Property, Associative Property of Addition, Combine Like Terms

OpenStudy (anonymous):

Thank you so much for posting the answer haitikidd001

OpenStudy (anonymous):

Yes thanks!! @haitikidd001